Space Stations

NASA International Space Station On-Orbit Status 15 August 2016

By Marc Boucher
Status Report
August 16, 2016
Filed under , ,
NASA International Space Station On-Orbit Status 15 August 2016
NASA International Space Station On-Orbit Status 15 August 2016.

Today: Mouse Epigenetics Cage Unit Maintenance: The crew completed standard maintenance for the Mouse Epigenetics experiment by refilling the Transportation Cage Units and Mouse Habitat Cage Unit with water and checking the water nozzles of the individual cages.
The Mouse Epigenetics investigation studies altered gene expression patterns in the organs of male mice that spend one month in space, and also examines changes in the deoxyribonucleic acid (DNA) of their offspring. Results from the investigation identify genetic alterations that happen after exposure to the microgravity environment of space.

NanoRacks Platforms-1 Module Removal: Four NanoRacks Modules were removed from NanoRacks Platform 1 and installed in the Minus Eighty-degree Freezer for ISS (MELFI). NanoRacks Modules 43 (Slime Mold), 44 (Awty-Yeast Cell Growth in a Microgravity Environment), 45 (Duchesne-Light Wavelengths on Algae Production), and 46 (Duchesne-Plant Growth Chamber) will remain in MELFI until they return on SpX-9.

At Home in Space Questionnaire: The CDR completed a questionnaire for the At Home in Space investigation. This Canadian Space Agency (CSA) experiment assesses culture, values, and psychosocial adaptation of astronauts to a space environment shared by multinational crews on long-duration missions. It is hypothesized that astronauts develop a shared space culture that is an adaptive strategy for handling cultural differences and they deal with the isolated confined environment of the space craft by creating a home in space.

NeuroMapping Operations: The crew performed testing in both a “strapped in” and “free floating” body configuration for this investigation which studies whether long-duration spaceflight causes changes to the brain, including brain structure and function, motor control, and multi-tasking abilities. It also measures how long it would take for the brain and body to recover from possible changes. Previous research and anecdotal evidence from astronauts suggests movement control and cognition can be affected in microgravity. The NeuroMapping investigation performs structural and functional magnetic resonance brain imaging (MRI and fMRI) to assess any changes that occur after spending months on the ISS.

Fluid Shifts Hardware Transfer and Service Module Setup: To prepare for Ultrasound activities in the Service Module (SM) this week, the crew transferred and set up hardware that supports the Fluid Shifts investigation from the Russian Segment. The experiment measures how much fluid shifts from the lower body to the upper body, in or out of cells and blood vessels, and determines the impact these shifts have on fluid pressure in the head, changes in vision and eye structures.

Fine Motor Skills: A series of interactive tasks on a touchscreen tablet was completed for the Fine Motor Skills investigation which tests skills needed to interact with technologies required in next-generation space vehicles, spacesuits, and habitats. The crewmember’s fine motor skills are also necessary for performing tasks in transit or on a planetary surface, such as information access, just-in-time training, subsystem maintenance, and medical treatment.

Extravehicular Activity (EVA) Preparations: The crew completed a procedures review in preparation for next Friday’s planned International Docking Adapter (IDA)2 EVA. Topics covered included prebreathe protocol review, Equipment Lock activities and suit donning plan, egress/ingress plan and EVA extension considerations. Following the review the crew participatet in a debrief with ground teams to discuss questions or concerns. The crew also verified that the Extravehicular Mobility Unit (EMU) glove heaters are functional and that the EMU TV is receiving power from the Rechargeable EVA Battery Assembly (REBA).

Waste and Hygiene Compartment (WHC): This morning the crew reported the WHC fan powered off during use which was followed by a “Check Separator Light” illumination. The crew performed the standard troubleshooting procedure to clear the Check Separator Light. The WHC was returned to normal use upon completion of the malfunction procedure. Additionally, the crew reported they have been seeing the “Check Separator Light” on previous use. The engineering team has scheduled a coordination meeting tomorrow to discuss these developments.

Mobile Servicing System (MSS) Mobile Transporter (MT) Translation: Today, ground controllers moved the MSS MT from work station 4 to workstation 6 in preparation for the removal of IDA2 from the SpaceX-9 trunk on Wednesday, and the installation of IDA2 onto PMA2 during the EVA on Friday.

LA Multiplexer/De-multiplexer (MDM) Patch – Ground Controllers successfully loaded a patch to the LA 1 MDM in preparation for IDA installation.

Payload MDM Transition – Ground Controllers were performing a scheduled PLMDM-2 High Rate Data Link (HRDL) reset which resulted in a complete lockup of the PLMDM-2 HRDL card. To recover from the HRDL card lockup, ground controllers performed a PLMDM transition. PLMDM-1 was subsequently powered ON and reconfigured as the primary PLMDM. All payload Health and Status (H&S) indications pre- and post-recovery were nominal. The PLMDM transition resulted in a loss of Health and Status for ~57 minutes.

Today’s Planned Activities
All were completed unless otherwise noted.

Install 17??.2076?-0 case onto SM ??? thermal control pipework r/g 3003
FINEMOTR – science ops run.
Video equipment installation for ??-2 exercise video shooting / r/g 2071
Replace MRM2;s [??1], [??2] dust filters and [?1], [B2] fan grids
On MCC Go test ??-2 ??-2 exercise – 3
MOUSE – h/w setup
DRAGON. Cargo transfer ops
WRS – water sample analysis
XF305 camcorder setup
EVA procedures’ final printout
MOUSE – module water replacement
MRM2 [??1] & [??2] air ducts vacuuming
Video equipment de-installation for ??-2 exercise video shooting r/g 2071
Mouse Epigenetics – Habitat 4 cleaning and Micro-G battery R&R, part 1
???’s [?-?] container and hose R&R. Post-R&R ??? activation / r/g 3061
DAN. Science Ops run r/g 2780
EVA procedure reviews
INTER-MAI-75. [???] h/w setup and activation. Refer to comment 6 r/g 3057
On-orbit life photography and video / r/g 2747
DAN. Science Ops run r/g 2780
Review EVA-related IDA installation procedure
AHIS. Questionnaire fill-up
FLUID SHIFTS. Equipment transfer to ROS.
FLUID SHIFTS. Ultrasound 2 hardware install in SM
Equipment preparation for PAO in LAB
FLUID SHIFTS. Ultrasound equipment connection to [???] / r/g 3054
Crew preparation for PAO / r/g 3060
PAO with Pyaty Element Project team r/g 3060
Mouse Epigenetics – Habitat 4 cleaning and Micro-G battery R&R, part 2
NMAP – science ops run
BIMS. Science Ops run r/g 3053
MRM2 Science Ops conference comm config
COULOMBIC CRYSTAL. H/w installation and setup r/g 3052
NANO. Module retrieval from the rack.
Regular preventive health care sprint exercise (at crew discretion).
DRAGON. Cargo transfer ops
COULOMBIC CRYSTAL. Science Ops run r/g 3052
INTER-MAI-75. H/w deactivation and closeout / r/g 3057
EDV (KOV) or EDV-SV fill (separation) for ELEKTROB. Inspect ??? No. 1212
NANO. Sample insertion into MELFI
Progress 432 (h/w comp’t) unstow and IMS update. Unstow ??? items from Progress 432 Zone 3
EVA procedure conference
??? maintenance
Return to MRM2 nominal comm config
COULOMBIC CRYSTAL. Data copy and downlink / r/g 3052
Glacier 2 trays removal
EVA camera D4 batteries charging
2.085 MSG file transfer
REBA batteries install
REBA-powered equipment c/o
Symbolic activity setup / r/g 3064
DRAGON. Cargo transfer ops
Downlink exercise data via ???
DRAGON. Dragon – cargo transfer conference
TOCA – data recording
HABIT reminder review
FLUID SHIFTS. Reminder 2 about Chibis suit measurements in SM

Completed Task List Items
Mouse M6 food R&R
EVA NZGL familiarization

Ground Activities
All activities were completed unless otherwise noted.
EVA prep
MT translate from WS4 to WS6
Nominal ground commanding

Three-Day Look Ahead:
Tuesday, 08/16: Fluid Shifts, FIR ACE T1 install, Microbial Air Sample pack for return, Mouse Epigenetics cage maintenance
Wednesday, 08/17: Fluid Shifts, OGA H2 sensor R&R, Dragon cargo ops
Thursday, 08/18: EVA procedures review/tool config/Equipment Lock prep, Mouse Epigenetics cage maintenance

QUICK ISS Status – Environmental Control Group:
Component – Status
Elektron – On
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Idle
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.